Case-Control Study:

Of 1200 workers in an asbestos factory, 55 people developed lung cancer. In a nearby factory of 4500 workers, 60 developed lung cancer. Consider asbestos as a cause of lung cancer.

1) Draw a 2x2 table & fill out the cells and label
2) Calculate prevalence of lung cancer among exposed
3) Calculate prevalence of lung cancer among non-exposed
4) Calculate prevalence of exposure among diseased
5) Calculate prevalence of exposure among non-diseased.
6) Calculate & interpret OR
Answer #1

Answer 1:

No of workers in asbestos factory: 1200

No of workers who developed lung cancer: 55

No of workers in nearby factory: 4500

No of workers who developed lung cancer: 60

Answer 2: Total number of workers in asbestos factory =1200

Number of workers who developed lung cancer = 55

Therefore, the prevalence of lung cancer among the exposed = 55/1200

                                                                                                 = 0.046

Answer 3: Total number of workers in nearby factory = 4500

Number of workers who developed lung cancer = 60

Therefore, the prevalence of lung cancer among the non-exposed = 60/4500

                                                                                                        = 0.013

Answer 4: Number of workers in both factories = 1200+4500 = 5700

Number of workers who developed lung cancer = 60+55 = 115

Therefore, the prevalence of exposure = 115/5700

                                                              = 0.02

Answer 5: Number of workers in both factories = 1200+4500 = 5700

Number of workers who did not develop lung cancer = 5700-115 = 5585

Therefore, the prevalence of non-exposure = 5585/5700

                                                                     = 0.98

Answer 6: OR (Odds Ratio)

                 = Odds of diseased in exposed group / Odds of diseased in non-exposed

= (55/1200)/(60/4500)

= 55/16

= 3.44

Therefore, the OR is 3.44 which is greater than 1. This means that the occurrence of lung cancer is more in the non-exposed group in compared to the exposed group.